The machine people are eager to turn over the machine to the exp, while continuing their development during the day hours. At todays 3 pm coordination meeting all exp said the machine was not ready for physics running, in particular the rf rebunching has only been done with few bunches in machine and with limite success (sigma ~40 cm, not the 20-25 required). Thus physics run will not start for another few days. Following this not all daytime hours will be for machine development, but which one is still undertemined. Though a shift schedule for the next week or so has been established, it it still subject to changes imposed by the machine and availability of people. Todays access was quite succesfull in identifying and fixing various problems. - BFS L1458 problem was solved changing mainframe. This implies there is no spare at this time. The DC and RICH should thus be able to take data in the coming days (an throughout the run) - All the si-wafers should now be operative. - several channels (TOFW, TFW2, BB) in terms of TDC's .. waere fixed by switching channels in tdc readout. - Magnet measurement of D1,D2 an D3 were done with calibrated Hall probes in order to solve some of the problems associated in particular with the high field measurements. The data still have to be digested, but it seems that significant differences in central field values from values deduced from the current are present at large filed settings. ---------------------------------------------------------------- Flemming Videbaek Physics Department Brookhaven National Laboratory e-mail: videbaek@bnl.gov phone: 631-344-4106 _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Mon Dec 29 18:31:44 2003
This archive was generated by hypermail 2.1.8 : Mon Dec 29 2003 - 18:32:06 EST